# Towards AI as Colleagues: Multi-Agent System Improves Structured Professional Ideation

### v1

@@ 3.1. Design Goals and Implementation
(1) DG1: Support adaptive Human-AI co-ideation dynamics. The system should enable users to fluidly navigate between exploratory and evaluative phases of collaborative ideation while maintaining strategic control over the creative process. This design goal was primarily motivated by prior research on the double diamond design process and human creative cognition patterns, which conceptualize innovation as an iterative progression through phases of divergent exploration and convergent refinement (Design Council, 2005).
(2) DG2: Enable rich, multi-perspective co-ideation. The system should facilitate engagement with diverse viewpoints and expertise domains to avoid narrow ideation patterns and encourage comprehensive exploration of solution spaces. This design goal was motivated by prior research on interdisciplinary collaboration and the limitations of single-agent interactions.
(3) DG3: Facilitate purposeful and transparent Human-AI collaborative control. To promote effective human–AI synergy, the interface should provide clear interaction points and user-friendly visual interfaces that enable users to maintain strategic oversight while leveraging AI capabilities for ideation support. This design goal emerged primarily from prior research on human-AI teaming, highlighting the need for balanced human agency and intuitive interface design in AI-mediated creative processes.
To address our design goals, we designed and developed MultiColleagues, a human–AI collaborative conversational platform where multiple AI personas participate in structured brainstorming alongside the user. The system architecture follows a two-tier design pattern: presentation layer (React frontend) and application layer (Flask API). The system integrates OpenAI GPT-4o for natural language generation (see Figure 2).
@@ 3.1.2. Multi-Persona Orchestration System
To enable rich, multi-perspective co-ideation (DG2), the system implements a persona orchestration framework that instantiates a diverse roster of AI colleagues with distinct professional backgrounds, communication styles, and domain expertise (see Figure 8 in Appendix D.1). Each persona is constructed through structured configurations defining behavioral instructions for communication patterns, specialized knowledge domains that establish topical authority, and participation patterns that govern engagement frequency. These persona templates are drawn from prior works on multi-agents’ interdisciplinary collaboration (Hong et al., 2024), which show that encoding standardized workflows into multi-agent prompts improves coordination and reduces cascading errors. We designed the personas to interact through a multi-stage selection process. First, all selected personas generate preliminary thoughts on the participant’s problem using individualized prompts. Then, an AI-driven algorithm evaluates these responses for relevance and engagement potential, selecting the first speaker to establish a natural conversation flow. For subsequent turns, a dynamic ranking mechanism selects the next speaking persona by scoring candidates on contextual relevance to prior user comments, conversation history, and unexpressed perspectives, while adding a 20% randomization factor to prevent deterministic patterns. The ranking algorithm instructs the language model to identify which persona would have “the strongest urge or most relevant comment to share next” given current discussion dynamics, ensuring diverse perspectives emerge organically rather than through artificial rotation. Once a persona is selected, the system retrieves its prompt template and combines it with the full conversation context, the user’s most recent input comments, and the current orchestration state (focus vs. explore). Guided by this structured input, the language model generates the persona’s output. To sustain longer dialogues, a conversational history compression pipeline is applied when the message count exceeds a threshold. As shown in Figure 9, recent turns’ history is kept in full while older persona contributions are summarized, allowing the system to maintain immediate context while compactly representing earlier perspectives. This compression design ensures efficiency and coherence in multi-colleague conversations (see Appendix D.2). Detailed prompt templates for persona creation, first-speaker selection, persona ranking, and response generation are provided in Appendix C.
3.1.3. User-Friendly Interface and Interaction Design
To maintain strategic oversight in multi-agent ideation while preventing passive consumption of AI output, we implement user-friendly interfaces and interaction mechanisms that enable clear control and integration points throughout collaborative discussions. Beyond the conversation flow controls described in DG1, the system provides clear visual cues through distinctive persona profile pictures and role-based message styling that enable users to quickly identify different AI perspectives and track individual contributions across extended discussions. An intelligent highlighting system with user-controlled visibility allows participants to manage information density by toggling keyword emphasis on or off, which supports cognitive load management and preserves access to AI-generated insights. The interface enforces each persona’s conversation wording limits and structured message threading that enhances readability during multi-agent exchanges. Besides, the thinking mode controls enable users to explicitly switch between “explore” mode (divergent thinking) and “focus” mode (convergent thinking), with clear visual indicators showing the current cognitive state and immediate effects on subsequent AI behavior.
Figure 3. Usage flow of the Multi-agent Conversational System for co-ideation. (1) Users first select a team of persona experts. (2) They then discuss the problem, with personas contributing from their own perspectives. The system supports two thinking modes: Explore mode expands ideas broadly, while Focus mode refines and synthesizes them. A summary recap highlights key points, and facilitation is triggered either by user request or automatically by the system when guidance is needed.:
3.2. Usage Scenario
4. Study Design
@@
Table 2. Participant demographics and creativity scores.
@@ 4.1. Participants
As shown in Table 2, our study recruited 20 participants through university mailing lists and professional networks (9 males, 11 females), aged 20-39 years (M = 26.7, SD = 3.8). 15 participants were undergraduate to PhD students pursuing degrees in fields such as computer science, information science, HCI, AI research, and communication, while the remaining 5 were early-career professionals working in technology, research, and design-related roles. All participants met our eligibility criteria of having backgrounds in relevant backgrounds and prior experience with creative problem-solving or AI-driven tools. This study received the university’s IRB approval. Each participant was compensated with \$20. Table 2 summarizes participants’ demographics along with their creativity scores (M = 5.39, SD = 1.16). Scores were calculated as the average of 11 items from a 7-point Likert-scale creativity assessment (see Appendix A.1 for the full questionnaire).
@@ 4.2. Study Procedure
Each study session lasted approximately 70 minutes and was conducted remotely via Zoom with a researcher present to observe interactions and provide technical support when necessary. All participants experienced both systems, MultiColleagues and ChatGPT (GPT-4o), in counterbalanced order to control for potential order effects (Pollatsek and Well, 1995). The detailed procedure comprised the following phases (see Figure 4 for visualized study workflow):
@@
Pre-study assessment (10 minutes): Participants completed demographic questionnaires capturing age, gender, educational background, and prior experience with AI-assisted creativity tools. Additionally, participants completed established creativity assessment instruments adapted from validated scales (Runco et al., 2001; Tierney and Farmer, 2002) to establish baseline creative capabilities (complete items provided in Appendix A.2).
@@
System interaction phase (20 minutes total; 10 minutes per system): Participants were randomly assigned to one of two condition orders, with equal distribution across sequences to control order effects. Think-aloud protocols (Van Someren et al., 1994) were employed throughout both interactions to capture real-time cognitive processes. For both conditions, a split-screen setup was provided with a Google Doc on the right side for recording generated ideas following each interaction. The allocated 10-minute timeframe could be extended if deemed insufficient, and early termination was permitted upon idea exhaustion or fatigue.
@@
Post-system evaluation (10 minutes total; 5 minutes per system): Following each system interaction, participants completed a 12-item survey on 7-point Likert scales (1 = strongly disagree, 7 = strongly agree). The survey evaluated three key dimensions: Experience (3 items), Outcomes (3 items), and System Design & User Control (4 items). For example, Outcomes evaluates if “I reached lots of valuable or actionable ideas that felt better than what I might have generated alone.” The complete set of 12 items is provided in Appendix A.
@@ 4.3. Evaluation Methods
To answer our research questions, we adopted a mixed-methods approach that integrates both behavioral and perceptual measures. Quantitative data were collected from conversation histories and system interaction logs, complemented by structured post-system evaluations of usability and creativity. Qualitative data came from semi-structured interviews, capturing participants’ experiences with collaboration strategies, the outcomes of navigating different interaction paradigms, and the utility of multi-agent versus single-agent approaches. The following sections detail our evaluation methods.
@@ 4.3.1. Comparative Analysis of User Ratings
To statistically evaluate the 12 survey data, we employed a structured analytical approach that began with thematically grouping the questions according to our research questions, followed by a non-parametric comparison of the two systems. The 12 questions were first organized into three core dimensions: (1) Experience (RQ1), which captured the subjective quality of the collaboration and user engagement, as well as factors shaping those outcomes; (2) Outcomes (RQ2), which measured the quality and novelty of the creative output; and (3) System Design & Control (RQ3), which evaluated the user’s perceived sense of control and flexibility. To streamline the analysis, closely related survey questions were merged into broader, more robust metrics, such as Outcome Quality & Novelty. Given the ordinal nature of the Likert scale data and the within-subjects design of the study, we used the non-parametric Wilcoxon signed-rank test to compare the paired ratings of MultiColleagues and Baseline for each metric. Effect sizes were calculated as the magnitude of observed differences between conditions.
@@ 4.3.2. Thematic Analysis
Two researchers conducted thematic analysis (Braun and Clarke, 2024) of interview transcripts, independently coding the data before reconciling differences and refining the codebook. Themes were organized around the study’s research questions on collaborative experience, creative outcomes, and system support, with sub-themes (e.g., “Traceable Perspectives”) emerging through iterative discussion and grouping into role- or condition-specific insights.
4.3.3. Topics Analysis and Conversational Structures
To complement engagement-level metrics, we examined the semantic structure of conversations by analyzing the number and distribution of discussion topics. We used GPT-5 with role-based prompts to segment transcripts into main topics (high-level themes) and sub-topics (supporting details). We chose GPT-5 as evaluator for its state-of-the-art reasoning and judgment capabilities across domains to ensure reliable idea assessment (Wang et al., 2025). To mitigate stochastic variability, each conversation was processed three times and the results averaged. All outputs were subsequently reviewed by a researcher for alignment with coding criteria. Based on these annotations, we calculated a branching ratio, defined as the mean number of sub-topics associated with each main topic. This metric characterizes whether conversations progressed more linearly (lower branching) or expanded into multi-threaded explorations (higher branching). To ensure comparability across sessions of different lengths, topic counts were normalized by conversation duration, producing topics per minute and sub-topics per minute as indicators of topical density. In addition, we calculated the average time per main topic and time per sub-topic to reflect the extent of conversational investment in individual ideas.
@@ 5. Results
We organize the results into two parts. First, we report log analyses that establish overall interaction statistics in the MultiColleagues condition (Section 5.1). Our descriptive results capture how participants engaged with the system in terms of colleague selection patterns and message distributions, providing context for interpreting subsequent findings. The remainder of the results sections (Section 5.1, 5.2, 5.3) are structured around our three research questions, combining survey responses, conversation logs, and interviews. For RQ1 (Experience), we examine how role-taking patterns and perceptions of social presence shaped participants’ collaborative atmosphere and engagement. For RQ2 (Outcomes), we assess how exposure to multiple AI-Colleague perspectives influenced the quality, novelty, and organization of creative outputs. For RQ3 (System Support), we evaluate how system design features affected participants’ sense of agency, control, and flexibility during ideation.
5.1. Overview of Participants’ Interaction Statistics
5.2. RQ1: Collaborative Experience with Multi-AI Colleagues
To address RQ1, we operationalized collaborative experience into three dimensions: complementary strengths, teammate-like relationships, and engagement. We observed significant differences across all, with MultiColleagues rated more positively than Baseline. We present results for each dimension below, drawing on both post-system evaluation responses (Table 3) and interviews.
@@ 5.2. RQ1: Collaborative Experience with Multi-AI Colleagues
5.2.2. Colleague Roles Enable Facilitative Leadership.
@@
5.2.3. Team-Like Atmospheres Shape Collaborative Experience.
To examine whether this heightened sense of team-like atmosphere was also reflected in participants’ own language, we analyzed linguistic cohesion metrics from Coh-Metrix (Graesser et al., 2004; McNamara et al., 2014) alongside pragmatic style ratings across conditions (see detailed methods in Appendix B.1). Results presented in Table 4 showed a significant difference in Directness, with the Baseline condition (M = 5.82, SD = 0.91) rated significantly higher than the MultiColleagues condition (M = 5.21, SD = 0.88; W = 30.0, $p$ = .009), suggesting that interactions with Baseline elicited more straightforward, task-focused language. Other contrasts did not reach statistical significance.
@@
5.2.4. Engagement and Flow Enhance Collaborative Immersion
Participants reported experiencing stronger engagement and conversational flow in MultiColleagues, with significantly higher ratings in the MultiColleagues condition (M = 5.70, SD = 1.38) compared to the Baseline condition (M = 4.45, SD = 1.57; W = 30.0, $p$ = .014). Behavioral interaction measures further reinforced this pattern (Table 5). Participants contributed nearly twice as many utterances in MultiColleagues (M = 8.35, SD = 5.79) than in the Baseline condition (M = 4.10, SD = 2.45; W = 9.5, $p$ = .001) and produced substantially more words overall (W = 18, $p$ ¡ .001). Sessions also lasted longer in MultiColleagues (M = 12.90, SD = 2.80) compared to Baseline (M = 9.80, SD = 2.30; W = 33, $p$ = .006), suggesting that participants remained more cognitively and temporally invested when coordinating multiple voices. Interview reflections echoed these dynamics. P5 explained, “I feel like in my discussions, I’m constantly prompting new information and asking for new information, kind of wanting to lead the discussion. I feel I’m more like a facilitator.” Others emphasized how the incremental delivery kept them attentive and invested (P3, P10, P12). While MultiColleagues fostered ongoing involvement, Baseline condition’s one-shot responses were often described as efficient but less engaging. P19 described the interaction as “relatively one-way, like a presentation… I just pick what I want,” and P8 admitting, “I didn’t feel like we were collaborating… I just wanted to hear its perspective.” These findings show that MultiColleagues’ structured rhythm sustained deeper immersion and engagement, while Baseline delivered efficiency at the cost of shallower participation.
5.3. RQ2: Impact of Multiple AI Perspectives on Creative Outcomes
To address RQ2, we examined participants’ evaluations of creative outcomes along three dimensions: creative exploration, process enrichment, and outcome quality and novelty. Results indicate that MultiColleagues was consistently rated more favorably than Baseline, with significant differences observed for creative exploration as well as for outcome quality and novelty (Table 6).

### v2

@@ 3.1. Design Goals and Implementation
(1) DG1: Support adaptive Human AI co-ideation dynamics. We focused on how co-ideation tools can help users shift smoothly between expansive exploration and targeted evaluation while maintaining control over the creative trajectory. Creative problem solving is understood as an iterative cycle of divergence and convergence that benefits from structured scaffolds (Design Council, 2005). Recent systems such as Supermind Ideator show that reflective moves and guided reframing can reduce fixation and improve idea quality (Heyman et al., 2024). Building on this work, our system supports flexible transitions across ideation phases as users’ goals evolve.
(2) DG2: Enable rich, multi-perspective co-ideation. Work on boundary objects demonstrates that shared artifacts help people coordinate across disciplinary differences by making reasoning and assumptions available for joint inspection (Star and Griesemer, 1989; Halpern et al., 2013). This aligns with creativity-support research showing that exposure to varied perspectives expands the idea space and reduces the risk of narrow, single-voice trajectories (Andolina et al., 2017; Siangliulue et al., 2016). More recent work with role-based and persona-based AI further shows that structured viewpoints can reveal tensions, question defaults, and surface assumptions that remain hidden when users interact with a single assistant (Pang et al., 2025; Zhang et al., 2024; Ghosh and Rintel, 2025; Liu et al., 2025). Building on this foundation, our system focuses on providing accessible and differentiated viewpoints that function as lightweight boundary objects, supporting interdisciplinary-style exploration even when users work alone and helping them avoid constrained or homogeneous idea paths.
(3) DG3: Facilitate purposeful and transparent Human-AI collaborative control. Human–AI teaming studies highlight that effective collaboration depends on shared awareness and clear communication that help calibrate trust between human and machine partners (Zhang et al., 2021, 2023). In creative and co-productive settings, coordination structures such as visible turn-taking, recognizable contribution types, and explicit interaction points help users interpret system actions and maintain oversight of the collaborative process (Winston and Magerko, 2017; Guzdial and Riedl, 2019; Kantosalo et al., 2020). Transparency and user control over automation are also linked to stronger engagement and more intentional steering of AI-generated content (Amershi et al., 2019; Zhang et al., 2025). Guided by this literature, our system emphasizes interaction patterns that make system behavior legible and predictable, offering clear interaction signals and visual scaffolds that allow users to work with AI colleagues while maintaining strategic control.
To address our design goals, we designed and developed MultiColleagues, a human–AI collaborative conversational platform where multiple AI personas participate in structured brainstorming alongside the user. The system architecture follows a two-tier design pattern: presentation layer (React frontend) and application layer (Flask API). The system integrates OpenAI Large Language Model GPT-4o for natural language generation (see Figure 2).
@@ 3.1.1. Adaptive Thinking Transition Mechanisms
Figure 3. Usage flow of the Multi-agent Conversational System for co-ideation. (1) Users first select a team of persona experts. (2) They then discuss the problem, with personas contributing from their own perspectives. The system supports two thinking modes: Explore mode expands ideas broadly, while Focus mode refines and synthesizes them. A summary recap highlights key points, and facilitation is triggered either by user request or automatically by the system when guidance is needed.
Composite UI showing team selection, problem discussion, Explore and Focus modes, summary recap panel, and facilitation panel.
@@ 3.1.2. Multi-Persona Orchestration System
To enable rich, multi-perspective co-ideation (DG2), the system implements a persona orchestration framework that instantiates a diverse roster of AI colleagues with distinct professional backgrounds, communication styles, and domain expertise (see Figure 7 in Appendix D.1). Each persona is constructed through structured configurations defining behavioral instructions for communication patterns, specialized knowledge domains that establish topical authority, and participation patterns that govern engagement frequency. These persona templates are drawn from prior works on multi-agents’ interdisciplinary collaboration (Hong et al., 2024), which show that encoding standardized workflows into multi-agent prompts improves coordination and reduces cascading errors. To support purposeful and transparent collaborative control, we implemented a structured turn-taking mechanism that manages how personas contribute during ideation (DG3). Personas introduce their perspectives in stages to reduce cognitive load and maintain contribution quality. Each persona first produces an individualized initial response, after which an AI-driven evaluation selects the opening speaker based on relevance and conversational coherence. For subsequent turns, a dynamic ranking mechanism identifies the persona most suited to speak next by assessing contextual relevance, conversation history, and unexpressed perspectives, with a small randomization factor (20%) to avoid rigid patterns. Once a persona is selected, the system retrieves its prompt template and combines it with the full conversation context, the user’s most recent input comments, and the current orchestration state (focus vs. explore). Guided by this structured input, the language model generates the persona’s output. To sustain longer dialogues, a conversational history compression pipeline is applied when the message count exceeds a threshold. As shown in Figure 8, recent turns’ history is kept in full while older persona contributions are summarized, allowing the system to maintain immediate context while compactly representing earlier perspectives. This compression design ensures efficiency and coherence in multi-colleague conversations (see Appendix D.2). Detailed prompt templates for persona creation, first-speaker selection, persona ranking, and response generation are provided in Appendix C.
Table 2. Participant demographics and creativity scores.
@@
3.1.3. User-Friendly Interface and Interaction Design
To maintain strategic oversight in multi-agent ideation while preventing passive consumption of AI output, we implement user-friendly interfaces and interaction mechanisms that enable clear control and integration points throughout collaborative discussions (DG3). Beyond the conversation flow controls described in DG1, the system provides clear visual cues through distinctive persona profile pictures and role-based message styling that enable users to quickly identify different AI perspectives and track individual contributions across extended discussions. An intelligent highlighting system with user-controlled visibility allows participants to manage information density by toggling keyword emphasis on or off, which supports cognitive load management and preserves access to AI-generated insights. The interface enforces each persona’s conversation wording limits and structured message threading that enhances readability during multi-agent exchanges. Besides, the thinking mode controls enable users to explicitly switch between “explore” mode (divergent thinking) and “focus” mode (convergent thinking), with clear visual indicators showing the current cognitive state and immediate effects on subsequent AI behavior.
3.2. Usage Scenario
4. Study Design
@@ 4.1. Participants
As shown in Table 2, our study recruited 20 participants through university mailing lists and professional networks (9 males, 11 females), aged 20-39 years (M = 26.7, SD = 3.8). 15 participants were undergraduate to PhD students pursuing degrees in fields such as computer science, information science, HCI, AI research, and communication, while the remaining 5 were early-career professionals working in technology, research, and design-related roles. All participants met our eligibility criteria, including relevant domain backgrounds and prior experience with creative problem-solving or AI-driven tools. The study received the university’s IRB approval, and each participant was compensated \$20. Table 2 summarizes participants’ demographics along with their creativity scores (M = 5.39, SD = 1.16). Scores were calculated as the average of 11 items from a 7-point Likert-scale creativity assessment (see Appendix A.1).
@@ 4.2. Study Procedure
Each study session lasted approximately 70 minutes and was conducted remotely via Zoom with a researcher present to observe interactions and provide technical support when needed. We used a within-subjects design to reduce individual variability in ideation skill and cognitive style, allowing direct comparison of collaboration patterns across conditions. All participants interacted with both MultiColleagues and ChatGPT, which shared the same underlying model GPT-4o, in counterbalanced order to mitigate order effects (Pollatsek and Well, 1995). ChatGPT was selected as the baseline as it represents the dominant single agent, chat based workflow for everyday ideation (Heyman et al., 2024; Shen et al., 2025; Wang et al., 2025b). This choice increases ecological validity by grounding the comparison in the AI as tool paradigm that users already rely on (Sinlapanuntakul and Zachry, 2025). While an ablated multi agent baseline could isolate specific mechanisms, our goal was not to test individual features in isolation but to examine whether a socially orchestrated AI as colleague system shifts how people brainstorm relative to this prevailing tool based model. The study procedure comprised the following phases (see Figure 4 for a visual overview):
@@
The figure outlines six sequential stages: introduction, pre-study assessment, problem formulation, system interaction, post-system evaluation, and semi-structured interview, which shows how participants progress through the user study.
@@
Pre-study assessment (10 minutes): Participants completed demographic questionnaires capturing age, gender, educational background, and prior experience with AI-assisted creativity tools. They also completed established creativity assessment instruments adapted from validated scales (Runco et al., 2001; Tierney and Farmer, 2002) to establish baseline creative capabilities (items provided in Appendix A.2).
@@
System interaction phase (20 minutes): Participants were randomly assigned to one of two condition orders, with equal distribution across sequences to control order effects. Think-aloud protocols (Van Someren et al., 1994) were employed throughout both interactions to capture real-time cognitive processes. For both conditions, a split-screen setup was provided with a Google Doc on the right side for recording generated ideas following each interaction. Each ideation session was set to approximately 10 minutes to emulate fast paced brainstorming, a duration informed by our pilot study which showed that the facilitator kept discussions on track and supported more focused topic oriented exchanges. This 10-min structure also aligns with prior work using short creative sprints for AI-assisted co-ideation (Pang et al., 2025; Qin et al., 2025). Our goal was to examine immediate coordination and convergence processes rather than long-term team formation. Following recent studies that allowed flexible engagement durations (Pang et al., 2025), participants could extend their sessions if additional time was needed or conclude early upon idea exhaustion or fatigue.
@@
Post-system evaluation (10 minutes total; 5 minutes per system): Following each system interaction, participants completed a 12-item survey on 7-point Likert scales (1 = strongly disagree, 7 = strongly agree). The instrument was adapted from established questionnaires (Brooke and others, 1996; Berretta et al., 2023; Carroll and Latulipe, 2009; Hart, 2006; Goodhue and Thompson, 1995). Items were selected from an initial pool of 30 and refined through iterative review for clarity and relevance to our research questions; a small pilot (N = 5) ensured comprehensibility and removed redundancy. The survey evaluated three key dimensions: Experience (3 items), Outcomes (3 items), and System Design & User Control (4 items). The complete set of 12 items is provided in Appendix A.
@@ 4.3. Evaluation Methods
To answer our research questions, we adopted a mixed-methods approach integrating behavioral and perceptual measures. Quantitative data were collected from conversation histories and system interaction logs, complemented by structured post-system evaluations of usability and creativity. Qualitative data came from semi-structured interviews, capturing participants’ experiences with collaboration strategies, the outcomes of navigating different interaction paradigms, and the utility of multi-agent versus single-agent approaches. The following sections detail our evaluation methods.
@@ 4.3.1. Comparative Analysis of User Ratings
To statistically evaluate the 12 survey data, we employed a structured analytical approach that began with thematically grouping the questions according to our research questions, followed by a non-parametric comparison of the two systems. The 12 questions were first organized into three core dimensions: (1) Experience (RQ1), which captured the subjective quality of the collaboration and user engagement, as well as factors shaping those outcomes; (2) Ideation Process and Perceived Outcomes (RQ2), which captured participants’ self-reported ideation process, their perceived outcome quality and novelty; and (3) System Design & Control (RQ3), which evaluated the perceived sense of control and flexibility. To streamline the analysis, closely related survey questions were merged into broader, more robust metrics, such as Outcome Quality & Novelty. Given the ordinal nature of the Likert scale data and the within-subjects design of the study, we used the non-parametric Wilcoxon signed-rank test to compare the paired ratings of MultiColleagues and Baseline for each metric. Effect sizes were calculated as the magnitude of observed differences between conditions.
@@ 4.3.2. Thematic Analysis
Two researchers conducted a thematic analysis (Braun and Clarke, 2024) of interview transcripts, independently coding the data before reconciling differences and refining the codebook. Themes were organized around the study’s research questions on collaborative experience, creative outcomes, and system support. Sub-themes (e.g., “Traceable Perspectives”) emerged through iterative discussion and were grouped into role- or condition-specific insights..
4.3.3. Analysis of User Idea Contribution
We examined the semantic structure of participants input during the ideation discussion to understand how they engaged with the task and how their ideas developed. GPT-5 with role based prompts was used to segment each user contribution into two levels. Main topics capture the central themes guiding the users reasoning. Sub topics capture the concrete points, reasoning paths, and directions expressed within those themes. We chose GPT-5 for its state-of-the-art reasoning and judgment capabilities across domains to ensure reliable idea assessment (Wang et al., 2025a). To reduce stochastic variation, each conversation was processed three times and the outputs were averaged. Two researchers then reviewed the extracted topics for alignment with the coding criteria, yielding strong inter rater reliability (Cohen’s $\kappa=0.86$) (Cohen, 1960). Based on these annotations, we computed a branching ratio defined as the number of sub topics divided by the number of main topics, which characterizes whether a participant developed ideas in a linear form with fewer branches or in a more exploratory form with multiple parallel directions. To further assess how participants expressed detailed concepts within their broader ideation problem, we applied a natural language toolkit for noun extraction in each user contribution (Bird et al., 2009). Nouns offer a reasonable proxy for concept introduction and have been used to estimate conceptual density in discourse analysis (Kintsch, 1998).
@@ 5. Results
We organize the results into two parts. First, we report log analyses that establish overall interaction statistics in the MultiColleagues condition (Section 5.1). Our descriptive results capture how participants engaged with the system in terms of colleague selection patterns and message distributions, providing context for interpreting subsequent findings. The remainder of the results sections are structured around our three research questions, combining survey responses, conversation logs, and interviews. For RQ1 (Experience), we examine how role-taking patterns and perceptions of social presence shaped participants’ collaborative atmosphere and engagement. For RQ2 (Outcomes), we assess how exposure to multiple AI-Colleague perspectives influenced the quality, novelty, and organization of creative outputs. For RQ3 (System Support), we evaluate how system design features affected participants’ sense of agency, control, and flexibility during ideation.
5.1. Overview of Participants’ Interaction Statistics
@@
5.2. RQ1: Collaborative Experience with Multi-AI Colleagues
To address RQ1, we operationalized collaborative experience into three dimensions: complementary strengths, teammate-like relationships, and engagement. We observed significant differences across all, with MultiColleagues rated more positively than Baseline. We present results for each dimension, drawing on both post-system quantitative survey evaluations (Table 3) and interviews.
5.2.2. Colleague Roles Enable Facilitative Leadership.
5.2.3. Team-Like Atmospheres Shape Collaborative Experience.
To examine whether this heightened sense of team-like atmosphere was also reflected in participants’ own language, we analyzed participants’ input through linguistic cohesion metrics from Coh-Metrix (Graesser et al., 2004; McNamara et al., 2014) alongside pragmatic style ratings across conditions (see detailed methods in Appendix B.1). Results presented in Table 4 showed a significant difference in Directness, with the Baseline condition (M = 5.82, SD = 0.91) rated significantly higher than the MultiColleagues condition (M = 5.21, SD = 0.88; W = 30.0, $p$ = .009), suggesting that interactions with Baseline elicited more straightforward, task-focused language. Other contrasts did not reach statistical significance.
Table 6. Statistical Comparison for Performance & Integration Metrics between MultiColleagues (MC) and Baseline. Results show significantly higher self-reported ratings for MC on Creative Exploration and Perceived Quality & Novelty.